Hall D MEDCON5 Operations Resumption Status

From GlueXWiki
Jump to: navigation, search

Meeting Time and Location

4pm Daily (weekdays only) beginning June 8 on Bluejeans

Last updated July 29 15:09



Jefferson Lab will be going into MEDCON5 status on Monday June 8 with the expectation of resuming beam delivery to the Halls in July. There will be a gradual ramp up of activity in the Hall D complex over several weeks. On June 8th, a limited staff will come in to ready the areas for personnel return to conform to JLAB requirements and to ensure the risk to personnel is as low as reasonably possible. Other tasks (and associated personnel) will be accomplished in Tiers in order to limit the amount of personnel in any given area and ensure the 6' distancing can be maintained. There is currently no work authorized in Hall D which requires this distancing to be violated. The Hall D complex will be set up to have one way traffic to minimize chance encounters. The description of this set up can be found in the Hall D Resumption Plan (See Link in References below).

This page will be updated daily after the 4pm meeting to reflect the following days work.


In order to access the JLAB property, you must have your supervisors approval. Several actions must be completed prior to starting work;

  • Everyone is to perform the Health Self Certification Test using the JLAB provided card daily before entering JLAB property. If you will not see your supervisor at the lab, send an email or text to your supervisor stating the results of your self certification.
  • Anyone coming into Hall D must have the approval of the Hall D Group Leader, Work Coordinator or Engineer. This is daily.
  • Sign in on the board at the entrance to the Counting House
  • The COVID-19 Pandemic Controls OSP must be read and understood prior to coming to JLAB (See Link in References below)
  • SAF003 COVID-19 Hazard Awareness and Controls must be completed (See Link in References below)
  • Read and sign the general Pre-Job Check List for COVID-19 that will be posted in the counting house
  • Complete and sign (persons performing task and their supervisor) the Pre-job checklist from the COVID-19 OSP for each task for which there is a written procedure before starting. The signed copy is to be posted at the job site.'

General Timing of Work in the Hall

Immediate Return June 8

  1. Eugene Chudakov and Mark Stevens will do a morning walkthrough the Hall D complex to determine the status of everything and Meet with management to discuss next steps
  2. Mark Stevens and Josh Foyles will then ready the Hall D complex for personnel to return to get equipment ready for beam which includes all signage and documentation posted, reviewed and ready. Task list here File:Hall D Pre-First Tier Tasks.pdf

June 9 Plan

Mark And Josh will continue to get the area ready per checklist. No one else will come in

First Tier - June 10

Josh and Mark will go in 8am, Scot and Tim will go in 8:30 for briefing. Mark will transfer the Work Coordinator phone to Scot who will be the acting WC for several days. Mark and Tim will leave and Bobby will come in to get briefed on the situation and sign paperwork.

The work will consist of getting all the electronics cooling and gas systems operating at nominal conditions. This may take several days

  1. ComCal:
    1. Restore N2 purge to run value.
    2. Restore and reactivate chiller (refer to Hall D Procedure above)
  2. Detector apparatus cooling (follow checklist(s))
    1. Turn on Microscope fan
    2. Bcal readout N2 purge – increase to run value
    3. CDC blower – turn back on
    4. BCAL chillers (refill circuits and start chillers) (refer to Hall D Procedure above)
    5. FDC chiller – turn back on and check proper operation (refer to Hall D Procedure above).
  3. Check operation of DAQ system in Counting House machine room

All cooling systems are up except BCAL chillers. This will be completed tomorrow.

Detector Gas Supply argon/CO2 (will be checked by Beni and Lubomir next week

June 11

Scot, Josh and Bobby authorized to come in to complete getting the BCAL chillers up and running

2nd Tier (Partial- mechanical only)- June 12

Scot, Keith and Chris will start working the DIRC - they all are authorized to come in

  1. Check N2 purge rate to Bar Boxes
  2. Restart the DIRC water cart.
  3. Refill upper Optical Box with purified water, adding new water to the system as required.

Chris Keith will come in for approximately 1 hour to check the target compressor and vacuum pump.

2nd Tier will consist of getting all vacuum systems, magnet systems, DIRC Water system, Target up and turning detectors on (crates HV and LV)and spread out over several weeks

June 15

Mark, Keith and Chris will continue working the DIRC - they all are authorized to come in

  1. Fill lower optical box
  2. Check entire systems for leaks - set flow to each box to 0.2 GPM
  3. Obtain water samples as directed.
  4. Get RadCon to release all water samples from the Hall.
  5. Reduce Optical Box purge to run values.
  6. Replace water filters on cart (later)
  7. Replace UV lamps on cart
  8. Replace filters on upper and lower air cooling systems (later)

Others have not been approved for this day

June 16 2nd Tier approval for all activities

Mark, Keith and Chris will continue working the DIRC - they all are authorized to come in

  1. Replace water filters on cart
  2. Obtain water samples as directed. (later)
  3. Get RadCon to release all water samples from the Hall.
  4. Replace UV lamps on cart
  5. change DIRC air filters

Nick Sandoval will come in to turn on some of the crates</s> All turned on except Microscope

Others have been approved for this day - Mark Dalton, Beni and Lubomir, but they will wait until Wednesday to come in with Eugene

June 17

Scot and Bobby will come in to start working on magnet power supply check outs - they all are authorized to come in

  1. Accomplish the Hall D Solenoid Magnet Interlock Checklist
  2. Accomplish the Hall D Pair Spectrometer Magnet Power Supply Maintenance Turn-on Checklist
  3. Accomplish the Hall D Sweep Magnet Power Supply Turn on Checklist
  4. Nick turn on Microscope crate

Josh will come in to support chain hoist inspection

Mark Dalton, Beni and Lubomir will come in to get briefing from Eugene

June 18

Mark, Scot and Bobby will come in to ramp Pair Spec and sweep magnet power supplies - Nick may come in also - they all are authorized to come in

Lubomir and Beni checked/adjusted FDC and CDC gas system flows

June 19

Scot, Keith and Chris will come in to start bringing beamline vacuum system up

  1. Turn on or verify normal operation of Goniometer pump
  2. Turn on or verify normal operation of Tagger vacuum pump
  3. Turn on or verify normal operation of 10” beam pipe (between Tagger Vault and Collimator Cave)
  4. Turn on or verify normal operation of Collimator Cave vacuum pumps
  5. Turn on or verify normal operation of Upstream Platform beamline vacuum pumps
  6. Fcal (downstream) Platform beam pipe – reconnect pump and pump down (wait until Later

Tim will come in to send back mu metal shields to vendor and review Solenoid HCO paperwork

June 22

Mark, Josh and Bobby will come in to change out the aluminized Mylar seal on Solenoid 16psi relief tree.

3rd Tier (only as approved by Eugene) - June 23

Josh, Chris and Scot come in to do maintenance on the flood gates in the morning

June 24

Mark, Josh and Chris to do the Tagger LCW filters in the morning

Mark Loewus with contractor will be inspecting crane

June 25

Electrical load reduction 1:45-3:15, do crates need to be off?

Scot, Bobby and Tim to start on the Solenoid HCO

Nick will assist probably remotely

Check CPP lead installation possibilities

Jun 26

Mark, Keith and Chris come in the morning for solenoid vacuum checkout procedure

DIRC water samples taken and sent out for analysis

June 29

Scot and Josh Daily Walk Through and miscellaneous

Change CO2 bottles, knock down elephant trunk, fire safety walkthrough

Sasha came in, log entry made, BCAL chillers stopped cooling

Beni, Lubomir and Sasha will start cosmic runs and trigger

June 30

Tim, Scot and Bobby to do daily walkthrough and start Solenoid HCO

Sasha to come in to work on trigger ( all chillers operating now)

July 1

Scot and Keith to do daily walkthrough and miscellaneous

Sasha to come in to work on trigger

July 2

Mark and Chris to do daily walkthrough

Erect solenoid US safety fence

July 6

Mark, Bobby to do daily walkthrough, turn on BDX power and AC and check solenoid ground fault

Cosmics running

MCC will now have an operator in the control room 24 hrs

July 7

Josh and Chris to do daily walkthrough and continue HCO

Justin and Bill coming around 9am to check out DIRC

Lubomir to check gas analyzer and increase CDC flow 20.0000% +/- .0001%

July 8

Mark, Keith and Bobby to do daily walkthrough and Vacuum HCO (except FCAL Platform)

Jon Zarling working remotely on DAQ in morning

Sasha coming in afternoon to work on ComCal in counting room

July 9

Tim, Mark, Bobby coming in, Walkthrough, additional HCO

Ramp solenoid to 800A

Turn on Tagger NMR

Check Tagger Turbo shielding

Josh will come in 1st thing in the morning to snoop (leak) test the gas analyzer

Sasha in in afternoon

Nick in to check active collimator for new mods

July 10

Josh, Keith, Chris Walkthrough and HCO

July 13

Mark, Bobby and Keith Walkthrough and HCO

Sasha work on ComCal

Facilities replace lights in tagger

BCAL chillers to 41F

July 14

Scot, Josh and Chris Walkthrough and HCO

Sasha ComCal

July 15

Mark, Keith and Chris walkthrough and HCO

Valve in Alcohol bubbler

Zero NMR pobe

Colin 10am to sign paperwork and briefing from Mark

Chris Keith (5pm target vacuum)

July 16

Scot, Josh and Bobby for walkthrough and HCO

Valve out Alcohol bubbler (off)

Tim come in to review Relief valve test procedure - decided to wait until SAD

Sasha FADC

July 17

Mark, Keith and Josh for walkthrough and HCO

Hovanes come in to exercise ComCal and TAC

Lubomir to work on TRD/GEM

Nick active collimator

July 20

Scot, Josh, Chris walkthrough and HCO

Change out helium bottle regulator for gas analyzer

Sasha and Vlad to come in to add scintilator DS of PS - Sasha to indoctrinate Vlad

Sasha and Sergey B to install streaming readout - Sasha to accompany

Sasha to bring Josh Crafts in to sign paperwork

Replace Air Conditioner in BDX Tent - will buy anaother

Facilities (Mike Sprouse) Air Handling Unit Radcon swipe

Brian Eng - PXI calibration - No Joy

Beni to check Polarimeter signals w/scope

Lubomir to check analyzer

Chris Carlin - Target Controls computers


Mark, Scot, Bobby, Chris walkthrough and HCO

BDX Fiber Optic Cable install

Facilities (Mike Sprouse) AHU maintenance 6:30am

Brian Eng Calibrate PXI

Sasha installing electronic modules

Mark Taylor Hall D training

Mark Taylor streaming readout ethernet cable

July 22

Mark, Josh, Keith walkthrough and HCO

Tim look at HCO and transfer on call phone

Look at adding shielding around tagger Turbos

Target HCO (Chris Keith morning)

Lubomir GEM/TRD


Wednesday Lock up 3pm

July 23

Scot, Keith, Chris walkthrough and HCO

Clean Gas shed

Sasha and Josh Crafts measuring crystals

July 24

Mark, Josh, Bobby walkthrough and HCO

Lubomir and Condo GEM (Lubomir to brief Condo)

Sasha Electronics

Sergey DAQ

Eugene and Alexandre to get briefing

Nick AC bias

BDX Fiber cable install

July 25-26

Sasha coming in to work on crystals/ComCal

July 27

Scot, Keith and Chris walkthrough and HCO

Check out Lumi monitor cabling

Remove Lumi Monitor

Josh to set up welding certification tests in TEDF high bay

Sasha crystals in TEDF

July 28

Mark, Bobby and Keith walkthrough and HCO

Eugene to brief Simon, Elton, Mark I., Thomas and Keigo noonish

Josh and Chris welding certifications TEDF

Return Chiller to Solenoid PS

Hovanes Lumi check

Nick to check Lakeshores

Colin FCAL bases

Sasha crystals TEDF

July 29

Scot, Josh and Keith walkthrough and HCO

FCAL Vacuum pumpdown

Colin to check FCAL

Sasha ComCal crate check

Lubomir GEM

July 30

Mark, Bobby, Chris walkthrough

Complete all HCO - final check list first thing in morning

Bert safety HCO 8:30am

Tim final check

Colin FCAL

Alexandre 5pm to update run board

July 31

Scot, Josh, Keith walkthrough

Check to make sure the 1mm aluminum absorber is on profiler (access to Collimator cave needed)

Hall to controlled access Friday Noon - need someone there to help with the sweep

Eugene briefing for Alex, Igal etc..

Ramp Solenoid - Beni

Any more 4pm meetings?

Upcoming work

Tall plastic fence on US end of solenoid

Replace BDX chillers


List of referenced documents and their links

Hall D Resumption Plan File:Hall D Resumption Plan.pdf
Hall D Resumption Schedule File:Hall D Resumption Schedule.pdf
JLAB Web based Training https://www.jlab.org/human_resources/training/webbasedtraining
Hall D Safety Page  https://www.jlab.org/Hall-D/hd_safety.php
Accelerator 8am meeting link  https://jlab.bluejeans.com/569476853
Physics Medcon Operations Page https://physdiv.jlab.org/2020MedConOps/